### v1

# LLaVA-UHD v2: an MLLM Integrating High-Resolution Feature Pyramid via Hierarchical Window Transformer

Abstract
In multimodal large language models (MLLMs), vision transformers (ViTs) are widely employed for visual encoding. However, their performance in solving universal MLLM tasks is not satisfactory. We attribute it to a lack of information from diverse visual levels, impeding alignment with the various semantic granularity required for language generation. To address this issue, we present LLaVA-UHD v2, an advanced MLLM centered around a Hierarchical window transformer that enables capturing diverse visual granularity by constructing and integrating a high-resolution feature pyramid. As a vision-language projector, Hiwin transformer comprises two primary modules: (i) an inverse feature pyramid, constructed by a ViT-derived feature up-sampling process utilizing high-frequency details from an image pyramid, and (ii) hierarchical window attention, focusing on a set of key sampling features within cross-scale windows to condense multi-level feature maps. Extensive experiments demonstrate that LLaVA-UHD v2 achieves superior performance over existing MLLMs on popular benchmarks. Notably, our design brings an average boost of 3.7% across 14 benchmarks compared with the baseline method, 9.3% on DocVQA for instance. We make all the data, model checkpoint, and code publicly available to facilitate future research.
1 Introduction
Figure 1: Comparison of LLaVA-UHD v2 with other MLLMs. (a) MLLMs typically align ViT features to language space using MLPs [63] or perceiver re-samplers [6, 52], lacking visual granularity. (b) Combining multiple visual encoders is non-universal and computationally intensive. (c) LLaVA-UHD v2 employs the Hiwin transformer to build an inverse feature pyramid and compress it into visual tokens, providing various semantic granularity for language generation.
Embedding visual information into large language models (LLMs) [97, 21, 4, 2] has significantly enhanced their ability to tackle complex visual multimodal tasks, such as visual question answering [8, 38, 32], document analysis [78, 77, 66], and visual interaction [19, 27]. With the rise of transformers [98], the vision transformer (ViT) [24, 85] has emerged as a standard approach for visual encoding in contemporary multimodal large language models (MLLMs) [23, 10, 99, 63, 62, 52, 4, 7, 103].
However, the ViT in an MLLM struggles with effectively handling all types of visual multimodal tasks, due to inadequate visual granularity captured by single-scale features, as illustrated in Fig. 1(a). For instance, visual grounding [107, 17, 110, 106] and optical character recognition [46, 49] demand fine-grained visual details, while image caption [84, 57] focuses on capturing high-level semantics and relations.
Drawing on the advancements in computer vision, comprehensive coverage of fine-grained details and high-level semantics can be integrated effectively by a feature pyramid [59, 87, 114]. However, this method presents two key challenges when applied to MLLMs: (1) Representation Inheritance. With its scalability, the ViT boasts general vision-language representations acquired through large-scale image-text pre-training ($e.g.$, CLIP [85] or SigLIP [109]), which is essential for continual learning in MLLMs. Modifying the ViT into a hierarchical architecture such as Swin [67] for multi-scale representations renders the vision-language representations of existing pre-trained models non-heritable. (2) Compression Effectiveness. The quadratic computational cost of LLMs with respect to the number of visual tokens necessitates effective compression of the feature pyramid. Current techniques [91, 60, 71, 96] often resize features to a fixed resolution, which distorts both the spatial shape and intrinsic semantics, as shown in Fig. 1(b). Moreover, down-sampling high-resolution features risks losing critical fine-grained details.
To address these challenges, we present LLaVA-UHD v2, an advanced MLLM centered around a novel Hierarchical window transformer, which enables capturing diverse visual granularities by constructing and integrating a high-resolution feature pyramid, as shown in Fig. 1(c). The Hiwin transformer performs an expansion-then-compression procedure on spatial resolution of feature maps, which consists of two stages: (i) Constructing an inverse feature pyramid. We train a parameterized feature up-sampler, the Joint Bilateral Upsampling (JBU) module proposed in FeatUp [26], on top of a pre-trained ViT ($i.e.$, CLIP-ViT [85]) to encode high-frequency information from the image into its feature map. This strategy can be applied to any ViT, inheriting its powerful feature representations, while constructing a multi-resolution feature pyramid based on the ViT. (ii) Integrating features with hierarchical window attention. To condense the multi-level feature maps effectively, we propose utilizing a set of hierarchical windows to capture semantics from local regions across different pyramid levels. A set of learnable queries is restricted to attend solely to key sampled features within their respective windows. This attention mechanism performs effective compression on local dense features at the native resolution of each pyramid level, thereby enabling visual tokens to effectively capture both fine-grained details and high-level semantics.
Extensive experiments demonstrate that LLaVA-UHD v2 dramatically outperforms the baseline method (LLaVA-UHD [31]) across 14 popular benchmarks, including document-centric visual question answering ($e.g.$, +9.3% on DocVQA), visual grounding ($e.g.$, average +5.7% on RefCOCOs [107]), and high-resolution image perception ($e.g.$, +3.4% on HR-Bench [100]). Moreover, we experimentally reveal that feature pyramids, regardless of the specific construction method ($e.g.$, bilinear interpolation), can further enhance the visual perception capabilities of MLLMs, offering new insights for future research. Our contributions can be concluded as follows:
  • We propose a visual pyramid encoding method by pre-training an up-sampling module, providing enriched semantic granularity for language decoding.
  • We propose a hierarchical window attention, capable of integrating and compressing a high-resolution feature pyramid into a condensed set of visual tokens, thereby ensuring computational efficiency while preserving the critical details needed for language decoding.
  • Trained efficiently on academic-scale data, LLaVA-UHD v2 achieves substantial improvements over the baseline method across 14 benchmarks.
2 Related Works
Figure 2: The overall architecture of proposed LLaVA-UHD v2, consisting of a ViT, our hierarchical window transformer (Hiwin transformer), and an LLM. Hiwin transformers process sliced patches and the overview image by capturing inner multi-level representations and compressing them into spatially consistent tokens for a better vision-language alignment.
2.1 Feature Pyramid Representation
Image pyramid techniques are fundamental in image processing, facilitating multi-resolution analysis since the era of manual feature design, as seen in SIFT’s scale-space keypoints [13, 70]. In deep learning, CNNs such as ResNet and VGG, inherently extract hierarchical features across scales [35, 93]. Innovations such as FPN and U-Net enhance semantic hierarchy for tasks like detection and segmentation [59, 87]. Recently, some transformer-based models [67, 111, 114] further advanced feature pyramid construction, capturing more comprehensive visual semantic granularity for visual representations. However, multimodal language models, often using CLIP-based ViTs, underutilize hierarchical features, suggesting a research gap for integrating advanced feature pyramids into these models [111, 85].
2.2 Visual Encoding in MLLMs
CLIP-ViT, favored for its effective alignment of visual features with linguistic semantics through contrastive pre-training, is widely adopted in MLLMs [85, 64, 63, 10, 62, 7, 113, 55]. Emerging research explores alternative visual representations, primarily in three categories: (1) Fusing features from CLIP-based CNNs and ViTs. LLaVA-HR [73] integrates stage-wise CNN features [68] into ViT’s layers, enhancing fine-grained perception. CogAgent [37] utilizes ViT features to query high-resolution CNN features for detailed information during language decoding. Mini-Gemini [54] employs a cross-attention-based post-fusion between CNN and ViT features. (2) Fusing features from visual experts trained with different vision pre-training tasks [60, 28, 71, 115, 96, 91, 102]. Candidate experts include DINO-v2 [83] by visual contrastive pre-training, SAM [47] by prompt segmentation pre-training, Pix2Struct [49] by document parsing pre-training, $etc.$ Deepseek-VL [71], SPHINX-X [60, 28] and Eagle [91] down-sample output feature maps and concatenate them along the channel axis. Cambrian-1 [96] initializes embeddings to query local patches from visual experts. (3) Language models for visual encoding: Fuyu [11], Otter-HD [51], and SOLO [20] encode images directly with LLMs, bypassing dedicated visual encoders. However, these approaches, while effective, increase computational demand and hinder a unified image-to-language design.
2.3 Token Projection and Compression
MLPs [64, 63], perceiver resamplers [6, 10, 103] and Q-Formers [52, 22, 112] are basic projectors widely used in modern MLLMs. Recently, various new designs have emerged. (1) Spatial-preserving compression. Qwen2-VL [99] and MiniGPT-v2 [16] employ a simple linear to merge tokens locally ($e.g.$ 2$\times$2). Honey-bee [15] introduces a C-abstractor ($i.e.$ CNN-based block), while Oryx [69] dynamically pools features and utilizes them to query the original ones. (2) Cross-layer feature compression. Token-Packer [53] compresses features across layers in a ViT using cross-attention. MMFuser [14] uses final-layer features as queries to attend to early-layer features. (3) Semantic-merged compression. Chat-UniVi [40] and LLaVA-PruMerge [89] extending the token merging [12] strategy, merge features with similar semantics into a single representation. However, these approaches depend heavily on feature padding, resizing, and reshaping, hindering their ability to compress features with arbitrary resolutions.
3 Method
3.1 Overview
The architecture of the proposed LLaVA-UHD v2 is illustrated in Fig. 2. It consists of three primary modules: a visual encoder (ViT), a vision-language projector (Hiwin transformer), and an LLM. We first divide and crop each input image into multiple slices with appropriate size and aspect ratio, leveraging the slicing strategy proposed in [31]. The image slices, along with an overview image, are then processed by a CLIP-ViT [85] with interpolated positional embeddings, capable of processing images with arbitrary sizes and shapes. The resulting features are subsequently passed to the Hiwin transformer for vision-language projection, which is carried out with two stages: (i) constructing an inverse feature pyramid (detailed in Sec.3.2), and (ii) integrating the feature pyramid by hierarchical window attention (detailed in Sec.3.3). The core of the Hiwin transformer lies in enhancing each ViT-encoded feature into a high-resolution feature pyramid encoding, thereby achieving enriched semantic granularity for each image slice. After this enhancement, visual tokens from different slices are reorganized into a spatially consistent feature map relative to the original image, ensuring clarity in spatial relationships. Following by concatenating with the overview tokens, all the visual tokens provide visual context for language decoding.
3.2 Inverse Feature Pyramid
Preliminaries. In the era of hand-crafted visual representations [70, 13], gaussian convolution filter is widely employed to construct image pyramids $\{\mathcal{I}^{l}\in\mathbb{R}^{\frac{H}{2^{l}}\times\frac{W}{2^{l}}\times 3},l=0,1...L-1\}$, where $l$ denotes the pyramid level, $(H,W)$ the image height and width, and $L$ the total number of levels. Building upon this concept, deep convolutional neural networks (CNNs), such as ResNet and VGG, naturally yield a bottom-up, hierarchical feature pyramid $\{\mathcal{F}^{l}\in\mathbb{R}^{\frac{H}{p\cdot 2^{l}}\times\frac{W}{p\cdot 2^{l}}\times C}\}$, where $C$ represents the feature dimension and $p$ denotes the a down-sampling ratio ($e.g.$, $p=8$ in ResNet-50 [35]). In such pyramids, lower-level feature maps generally have higher resolution and capture more visual detail, whereas higher-level feature maps exhibit lower resolution but contain more abstract semantic information. However, ViTs splitting the image into coarse patches, only produce single-scale feature maps ($i.e.$, $\mathcal{F}^{0}\in\mathbb{R}^{\frac{H}{p}\times\frac{W}{p}\times C}$, where for instance, $p=14$), as illustrated in Fig. 1(a). Lacking such feature pyramid, like that present in CNNs, hinders the performance of ViTs on MLLM tasks, which require both fine-grained visual details and high-level semantic information. Consequently, how to construct a ViT-based feature pyramid with varying semantic granularity remains an unresolved challenge.
Figure 3: Flowchart of the Joint Bilateral Upsampling (JBU) module, which leverages the image pyramid to guide feature up-sampling, integrating high-frequency information into the up-sampled feature maps.
Guided feature up-sampling. Due to the lack of high-resolution features, up-sampling ViT features becomes the necessary strategy to inversely construct the feature pyramid. Two straightforward approaches can be employed:(1) plain bilinear interpolation and (2) deconvolution network. By doubling and quadrupling the resolution of resulting feature maps, a ViT-based feature pyramid $\{\mathcal{F}^{l}\in\mathbb{R}^{\frac{H\cdot 2^{l}}{p}\times\frac{W\cdot 2^{l}}{p}\times C},l=0,1,2\}$ is constructed. While effective, directly up-sampling deep features hardly introduces precise visual details, which are essential for tasks requiring fine-grained visual information. To address this limitation, we leverage the parameterized JBU module [26] to guide feature up-sampling with image priors, which helps introduce high-frequency visual details.
Generally, the objective of JBU module is to learn $(l-1)$ convolutional layers on the image pyramid $\{\mathcal{I}^{l}\in\mathbb{R}^{\frac{H\cdot 2^{l}}{p}\times\frac{W\cdot 2^{l}}{p}\times 3}\}$ to capture high-frequency patterns of local texture for guiding feature up-sampling, in Fig. 3. Specifically, for each input image, its $(l+1)$-th level ViT features is defined as
$\begin{split}\mathcal{F}^{l+1}=\mathrm{Conv}\big{(}\mathrm{Up}(\mathcal{F}^{l});\Theta^{l+1}(\mathcal{I}^{l+1})\big{)},\end{split}$ (1)
where $\mathrm{Up}(\cdot)$ denotes the up-sampling interpolation and $\mathrm{Conv}(\cdot)$ represents the convolutional operation applied to feature maps using customized kernel weights $\Theta^{l+1}$ learned from the corresponding image $\mathcal{I}^{l+1}$.
As a result, each pixel value of the up-sampled feature maps is extracted as
$\mathcal{F}^{l+1}[x,y]=\frac{1}{|U|}\sum_{(x^{\prime},y^{\prime})\in U}\biggl{(}\mathrm{Up}({\mathcal{F}^{l}})[x^{\prime},y^{\prime}]\times D_{dist}\times D_{sim}\big{(}\Theta^{l+1}(\mathcal{I}^{l+1}[x,y]),\Theta^{l+1}(\mathcal{I}^{l+1}[x^{\prime},y^{\prime}])\big{)}\biggr{)},$
where $[x,y]$ denotes the coordinate index of feature maps, $U$ the region of convolution neighborhood, $D_{dist}$ is a decay factor based on the distance between $[x,y]\ \text{and}\ [x^{\prime},y^{\prime}]$, and $D_{sim}$ is a similarity weight based on the attention of each image pixel using a two-layer GeLU MLP $\Theta^{l+1}$. Additional details can be found in supplement A.1. By iteratively performing Eq. 1, we progressively up-sample the feature maps and ultimately construct the feature pyramid $\{{\mathcal{F}^{l},l=0,1,2}\}$.
Learnable up-sampler optimizing. Following [26], we optimize the reconstruction loss between the highest-level feature maps and the lowest as
$\mathcal{L}=\lVert\mathcal{F}^{0}-\mathrm{Down}(\mathcal{F}^{2};\Omega)\rVert_{2}^{2},$ (3)
where $\mathrm{Down}(\cdot)$ is a down-sampling network with trainable weights $\Omega$, $e.g.$, two convolution layers with a stride of 2. Note that we omit some hyper-parameters in Eq. 3 for clarity. Please kindly refer to the supplement A.1 for details.
However, when training the JBU module upon CLIP-ViT [85], we encounter the texture detail degradation of high-resolution feature maps. Specifically, features of CLIP-ViT scarcely exhibit enhanced detail representation at higher resolutions, compared to other ViTs like DINO-ViT [83]. We attribute this issue to the lack of inter-level feature reconstruction for language-aligned features during iterative up-sampling, resulting in an unstable transformation of such semantic-centric representations of the CLIP-ViT. To address this, we propose a straightforward hierarchical supervision strategy, regulating each level in the pyramid. Eq. 3 is updated as follows. The complete form for Eq. 3 and additional details can be found in supplement A.1.
$\mathcal{L}=\frac{1}{2}\sum_{l=1}^{2}\lVert\mathcal{F}^{0}-\mathrm{Down}(\mathcal{F}^{l};\Omega^{l})\rVert_{2}^{2}$ (4)
Figure 4: The flowchart of hierarchical window attention. Feature maps from different levels of the feature pyramid are adaptively RoI-aligned into sampling features and then concatenated along the length axis to serve as the key for the learnable queries.
3.3 Hierarchical Window Attention
The hierarchical nature of feature pyramids necessitates an effective approach for compressing features at varying resolutions while maintaining cross-level spatial alignment.
Hierarchical window generation. Inspired by object detection [86, 114], we employ the RoI-align [36] to sample key features while preserving the spatial locality across feature maps at different levels, as illustrated in Fig. 4. Specifically, we begin by uniformly dividing the feature maps of each level into $N\times N$ windows, whose widths and heights are float-point values $(\frac{W}{N},\frac{H}{N})$ rather than integers. Especially, windows that share the same “anchor” points across levels form a set of hierarchical bounding boxes, represented by the top-left and bottom-right coordinates $\{\mathcal{R}^{l}_{i,j}\in\mathbb{R}^{1\times 4},i,j\in 0,1,2...N-1,l=0,1,2\}$, where $l$ indicates the feature level and $(i,j)$ the 2D indices. To mitigate the size distortion of feature maps caused by the discrepancy between the aspect ratio of RoI-aligned feature maps and the original feature maps, we define a pooling score to evaluate this discrepancy:
$\begin{split}S(W,H,r_{w},r_{h})=-\left|\log\frac{W}{H}-\log\frac{r_{w}}{r_{h}}\right|,\end{split}$ (5)
where $(r_{w},r_{h})$ denotes the width and height of pooled features. By maximizing the score $S$, we select the optimal grid size $(r^{*}_{w},r^{*}_{h})$ from a set of predefined proposals $\{(3,3),(2,3),(3,2),(2,4),(4,2)\}$. Subsequently, we perform RoI-align with the generated windows to sample the key feature maps for the following attention operation.
Cross-window feature querying. To compress the multi-level feature pyramid $\{\mathcal{F}^{l},l=0,1,2\}$ of a given image slice $\mathcal{I}$, we initialize a set of queries $\{\mathcal{Q}_{i,j}\in\mathbb{R}^{1\times C},i,j\in 0,1,2...N-1\}$, each of which corresponds to a set of hierarchical windows $\{\mathcal{R}^{l}_{i,j},l=0,1,2\}$, in Fig. 4. Given each query vector $\mathcal{Q}_{i,j}$, we prepare the key vector $\mathcal{K}^{l}_{i,j}\in\mathbb{R}^{(r^{*}_{x}\cdot r^{*}_{y})\times C}$ of $l$-th level as
$\begin{split}\mathcal{K}^{l}_{i,j}=\text{RoI}(\mathcal{F}^{l},\mathcal{R}^{l}_{i,j})+\phi^{l},\end{split}$ (6)
where $\phi^{l}$ is a level-wise positional embedding. The key vectors $\mathcal{K}^{l}_{i,j}$ are then concatenated along the length axis to form the final key vector $\mathcal{K}_{i,j}\in\mathbb{R}^{(3\cdot r^{*}_{x}\cdot r^{*}_{y})\times C}$ for each query $\mathcal{Q}_{i,j}$. The corresponding value vector $\mathcal{V}_{i,j}$ is obtained in the same manner, but without the level-wise positional embedding. Thus, the cross-attention can be performed as
$\mathcal{Q^{*}}_{i,j}=\text{CrossAttn}(\mathcal{Q}_{i,j}+\varphi_{i,j},\mathcal{K}_{i,j}+\phi_{i,j},\mathcal{V}_{i,j}),$ (7)
where $\mathcal{Q^{*}}_{i,j}$ denotes the updated query, and $\varphi,\phi$ is the 2D spatial position embedding of query and key vector, respectively. In the end, we concatenate all the updated query $\mathcal{Q^{*}}_{i,j}$ into a feature map $\mathcal{P}\in\mathbb{R}^{N\times N\times C}$ to represent the visual token of image $\mathcal{I}$.
3.4 Spatially-consistent Token Organization
Due to the varying partitions across different images, effectively organizing and conveying the structure of slices to the LLM is crucial for a more accurate understanding of the image. Prior studies [31, 60, 103] employ spatial tokens (e.g., “\$\backslash n$” or “,”) to denote the relative positioning of image slices. While such approaches provide cues regarding the global arrangement of each slice, they overlook the intrinsic spatial relationships among individual visual tokens. For instance, two horizontally adjacent tokens in the image, located in different slices, may become significantly separated in a 1D arrangement. To address this, we leverage our Hiwin transformer, which preserves 2D spatial consistency with the original image, and amalgamate these 2D feature maps into a large 2D feature map according to the slicing configuration, in Fig. 2. A newline token “\$\backslash n$” is then inserted between rows, followed by flattening them into the 1D arrangement. Additionally, a comma “,” is placed between the large feature maps and the features of the overview image to distinguish between different views of the image.
4 Experiment
In this section, we present an empirical evaluation of LLaVA-UHD v2. We begin with a comprehensive outline of the implementation details of our model, followed by a comparative analysis of its performance across widely recognized benchmarks against competitive counterparts. Finally, we provide an in-depth ablation to further elucidate the capabilities and behavior of LLaVA-UHD v2.
4.1 Implementation Details
Model Setting. We adopt LLaVA-UHD [31] as the baseline method. Specifically, we employ CLIP-ViT-L/14-336 as the visual encoder, Vicuna-7B [21] as the language model, and our proposed Hiwin Transformer as the vision-language projector. We set the maximum slice number to 6 to cover a range of aspect ratios and image resolutions. The number of learnable local queries is set as 144 ($i.e.,$$N=12$). LLaVA-UHD v2 consists of a three-stage training process, as outlined below.
Stage 0: JBU module pre-training. Before employing the JBU module in MLLM, we pre-train it on the MS-COCO [58] dataset with a global batch of 16 on 8$\times$×A100. We leverage Adam optimizer with a learning rate of 1e-3 for 2000 steps. During this stage, the parameters of CLIP-ViT are frozen.
Stage 1: MLLM pre-training. In this stage, the parameters of the visual encoder, JBU module, and LLM are frozen. We only fine-tune the parameters within the hierarchical window attention of the HiWin Transformer using LLaVA-Pretrain [63] dataset for 1 epoch. We employ the AdamW optimizer with a learning rate of 1e-3 and a cosine learning rate schedule. The global batch size is set to 256. Note that during this stage, we only encode the overview image, excluding the slices, for efficiency.
[…]
4.2 Experimental Setting
We present the experimental settings, detailing the benchmarks, evaluation metrics, and compared counterparts.
Benchmarks. Extensive benchmarks are used to analyze the effectiveness of our modules. We categorize these benchmarks into the following groups: (1) General VQA benchmarks including MME [25], MMB [65], SEED-Image [50] and GQA [38]; (2) Knowledge-based VQA benchmarks including MMMU-val [108], Science-QA [72], AI2D [44]; (3) OCR-based VQA benchmarks including ChartQA [77], OCR-Bench [66], TextVQA [94] and DocVQA [78]; (4) Visual spatial understanding benchmarks such as RealWorldQA [1] and RefCOCOs [107] (used in ablation studies); (5) High-resolution image perception benchmarks like HR-Bench(4K) [100].
Evaluation Protocols. Beyond benchmark evaluations, we report additional metrics for comprehensive analysis, including: (1) overall volume of training data, (2) maximum supported image resolution for each method, and (3) computational cost of the entire MLLM at maximum resolution.
Counterparts. We compare our model with the advanced MLLM counterparts. (1) General MLLMs like Qwen-VL [10], MiniGPT-v2 [16], Honey-bee [15], mPLUG-Owl2 [105], VILA [56] and LLaVA-1.5 [63]. (2) High-resolution MLLMs including UReader [104], Monkey [55], LLaVA-HR [73] ,LLaVA-Next [62]. (3) Mixture of visual experts such as  [73], SPHINX-series [60, 28] and Mini-Gemini [7].
Table 1: Main performance on popular benchmarks. For a fair comparison, we only report the method using 7B level LLM ($e.g.$, Vicuna-7B). #Data denotes the volume of overall data during MLLM pre-training and supervised fine-tuning. “MaxRes.” is the maximum accessible resolution of MLLM. “Avg.”: average results of 13 benchmarks. “VQAD: DocVQA. “BenchOCR”: OCR-Bench. “VQAC”: ChartQA. “VQAT”: TextVQA. “SQA”: Science-QA. “MMMUv”: MMMU-val. “SEEDI”: SEED-Image. “MMEP”: perception sub-set of MME. “RWQA”: RealWorldQA. “BenchHR”: HR-Bench.
OCR & Chart Knowledge General Vision Spatial High Res.
Method #Data MaxRes. #FLOPs. Avg. VQAD BenchOCR VQAC VQAT AI2D SQA MMMUv GQA SEEDI MMB MMEP RWQA BenchHR
Qwen-VL [10] 1.45B 448$\times$448 4.0T 56.9 62.6 48.8 66.3 61.5 57.7 68.2 35.9 57.5 65.4 61.8 74.4 49.3 30.5
MiniGPT-v2 [16] 326M 448$\times$448 4.0T - - 15.7 - - - - - 60.3 - - - - -
mPLUG-Owl2 [105] 401M 448$\times$448 1.7T - - - - 58.2 - 68.7 - 56.1 57.8 64.5 72.5 - -
UReader [104] 86M 896$\times$1120 20.3T - 65.4 - 59.3 57.6 - - - - - - - - -
LLaVA-1.5 [63] 1.22M 336$\times$336 8.0T 49.0 21.8 31.8 17.8 45.5 55.5 66.8 37.0 62.0 65.8 66.5 75.3 54.8 36.1
SPHINX-2k [60] 1.01B 762$\times$762 42.2T - - - - 61.2 - 70.6 - 63.1 71.6 65.9 73.6 - -
SPHINX-X [28] 15.3M 448$\times$448 21.3T - 56.3 - 39.7 58.1 63.0 70.4 - 56.2 68.8 57.9 63.0 - -
LLaVA-HR [73] 1.22M 1024$\times$1024 24.3T - - - - 67.1 - 65.1 - 64.2 64.2 - 77.7 - -
VILA [56] 51M 336$\times$336 8.2T - - - - 64.4 - 68.2 - 62.3 61.1 68.9 76.7 - -
Honey-bee [15] 52.5M 336$\times$336 2.6T - - - - - - - 35.3 - 64.5 70.1 79.2 - -
Mini-Gemini [54] 3.0M 672$\times$672 54.6T 59.4 61.9 47.7 47.4 65.2 68.2 69.6 36.8 64.5 66.9 65.8 77.3 51.1 50.1
Monkey [55] 1.40B 896$\times$1344 28.0T 59.2 66.5 51.4 65.1 67.6 62.6 69.4 38.9 60.7 64.3 59.8 73.6 51.6 38.0
LLaVA-Next [62] 1.34M 672$\times$672 44.4T 61.0 63.6 53.2 54.3 64.9 67.0 70.1 35.8 64.2 70.2 67.4 76.0 57.8 47.9
LLaVA-UHD v2 (ours) 1.42M 1008$\times$672 17.5T 63.2 68.1 53.9 64.5 67.6 70.5 71.3 38.2 65.4 70.0 68.2 74.7 58.2 51.5
Table 2: Ablation studies of modules in our proposed method. “HFP” is the abbreviation of high-resolution feature pyramid. “$\Delta$” denotes the overall improvement compared to the baseline. REC reports the average accuracy of RefCOCO/g/+.
OCR & Chart Knowledge General Vision Spatial High Res.
Method Average VQAD BenchOCR VQAC VQAT AI2D SQA MMMUv GQA SEEDI MMB MMEP RWQA REC BenchHR
LLaVA-UHD [31] 58.0 56.7 40.9 56.3 62.2 55.4 70.7 37.0 63.8 65.6 64.8 70.0 54.4 68.3 45.6
$+$ JBU module 60.0 60.2 50.4 60.4 67.1 57.8 70.5 38.2 64.0 66.7 65.6 71.2 51.9 72.3 43.9
$+$HFP integration 61.5 65.0 51.3 62.5 68.5 58.1 69.2 38.9 64.6 67.4 65.5 73.0 55.5 73.3 48.9
$+$Token organization 61.7 66.0 50.1 62.8 66.8 59.4 69.8 37.6 64.0 67.4 66.1 73.6 56.9 74.0 49.0
$\Delta$ +3.7 +9.3 +9.2 +6.5 +4.6 +4.0 -0.9 +0.6 +0.2 +1.8 +1.3 +3.6 +2.5 +5.7 +3.4
4.3 Main Performance
Table 1 showcases a comparative analysis of our proposed LLaVA-UHD v2 against state-of-the-art MLLMs across 13 widely recognized benchmarks. To facilitate a comprehensive evaluation, we have selected 5 prominent open-source models, and assessed them on all 13 benchmarks, to serve as our reference points. (1) LLaVA-UHD v2 outperforms current counterparts. Our model exhibits a significant enhancement in average performance, achieving an improvement of 2.2% compared to LLaVA-Next, among other models. Specifically, compared with general models (such as Qwen-VL and LLaVA-1.5) and high-resolution MLLMs (like Monkey and LLaVA-Next), LLaVA-UHD v2 demonstrates consistent improvements across various tasks, including general VQA (e.g., 65.4% on GQA), ultra-high-resolution image perception (e.g., 51.5% on HR-Bench). Notably, LLaVA-UHD v2 surpasses OCR-centric models like UReader on DocVQA (68.1% vs. 65.4%) and outperforms those with multiple experts (such as SPHINX-2k and Mini-Gemini), achieving superior performance on general tasks like MMB (68.2%) and SEED (70.0%). These results highlight the value of rich semantics derived from multi-level visual granularity in enhancing both the understanding and perceptual capabilities of MLLMs. (2) LLaVA-UHD v2 indicates efficiency on data utilization and computation. Compared to LLaVA-Next and Mini-Gemini, both operating at the 672$\times$672 resolution, LLaVA-UHD v2 supports 1.5 times the resolution ($i.e$., 672$\times$1008) and achieves superior performance with less than 40% of the computational cost. Furthermore, in contrast to Honey-bee and VILA, which utilize 52.5M and 51M data samples respectively, LLaVA-UHD v2 attains comparable or superior performance using only $\sim$2.8of the data, demonstrating the data efficiency of our model, making it highly suitable for low-cost exploratory research in the academic community.
Figure 5: Performance on different visual tasks with JBU module and vanilla bilinear interpolation. “OCR” denotes the optical character recognition, “Seg” the Linear probing semantic segmentation, and “Cls” the fine-grained classification on SUB-200.
Table 3: Comparison of different methods for feature pyramid construction. “ConvNext” means we replace the CILP-ViT with CLIP-ConvNext [68] as visual encoder and directly use the feature maps from multiple stages as the final hierarchical feature pyramid.
General Knowledge OCR & Chart High Res.
Method Average MMEP GQA AI2D VQAC VQAT VQAD BenchHR
LLaVA-UHD 58.6 70.0 63.8 55.4 56.3 62.2 56.7 45.6
w. ConvNext 59.7 68.2 62.7 55.6 61.8 63.5 61.8 44.0
w. DeConv. 61.7 71.2 64.2 57.4 61.8 67.8 63.4 46.3
w. Bilinear 62.0 72.0 64.5 57.8 62.2 67.6 63.7 46.5
w. JBU module 63.0 73.0 64.6 58.3 62.5 68.5 65.0 48.9
4.4 Analytical Study
We conduct analytical experiments on proposed modules to validate the effectiveness of LLaVA-UHD v2. Note that we freeze the CLIP-ViT during the supervised fine-tuning, which facilitates a clearer verification of the effectiveness of each module introduced in the Hiwin transformer.
Main module ablation. In Table 2, by replacing the low-resolution CLIP-ViT features with JBU-enhanced high-resolution ones, 2% average improvement can be seen, especially, on tasks depending on visual details like OCR-Bench (+9.5%) and RefCOCOs (+4.0%). Employing the Hiwin transformer to integrate the feature pyramid constructed by the JBU module further improves average accuracy by 1.5%, especially 5.0% on ultra-high resolution perception (HR-Bench), demonstrating rich visual granularity facilitates more precise language generation. Additionally, spatially consistent token organization leads to further performance improvements in visual-spatial understanding, with a 1.4% increase on RealWorldQA and a 0.7% improvement on RefCOCOs.
Figure 6: Qualitative comparison of proposed LLaVA-UHD v2 and advanced MLLMs, including LLaVA-Next, Mini-Gemini, and GPT-4V on high-resolution complex perception tasks, which require the integration of both fine-grained visual information and high-level semantic contexts.
Table 4: Comparison of different choice of grid sizes on performance and efficiency. “Pyramid” means the feature grids from different levels form a region-level feature pyramid, $e.g.$, [2$\times$3] for level-0, [4$\times$6] for level-1, [8$\times$12] for leval-2. “Fix” represents all feature maps are pooled into a 3$\times$3 feature grid. We measure the training period on 8$\times$×A100s, the latency on an A100 with a 1008$\times$672 image, and the GPU memory on 8$\times$×A100s with 1 image per GPU in supervised fine-tune phase.
Efficiency General Knowledge OCR & Chart
Method Period(h) Latency(s) Memory(G) Average MMEP GQA AI2D VQAC VQAT VQAD
Pyramid 62.4 1.26 60.3 62.4 69.0 60.8 57.3 60.7 67.5 58.9
Fix [3$\times$3] 26.9 0.62 41.7 64.6 73.8 63.9 58.8 60.9 66.2 63.8
Selective 27.7 0.54 39.4 65.3 73.0 64.6 58.3 62.5 68.5 65.0
JBU module demonstrates effectiveness on both MLLM tasks and vision tasks. We investigate the generalizability of different feature up-sampling methods ($i.e.$, bilinear interpolation, de-convolution [36], JBU module, and CNN [68]) across diverse tasks. (1) Results on MLLM tasks. Table. 3 clearly shows that feature pyramids, regardless of their construction method, can enhance performance across various MLLM tasks. Nonetheless, the feature pyramid constructed with our JBU module achieves an average performance gain of 1.0% over bilinear interpolation, indicating that the JBU module further enhances beneficial visual representations ($e.g.$, high-frequency visual features). (2) Results on visual tasks. Beyond the VQA task in MLLM, we further select some fundamental visual tasks, including semantic segmentation [58], optical character recognition [80], and fine-grained classification [45], to compare the effect of bilinear interpolation and JBU module. The settings of training and testing are detailed in supplement A.2. As shown in Fig. 5, the JBU module outperforms the bilinear interpolation on OCR (+4.0%), semantic segmentation (+4.7%) and fine-grained classification (+3.8%), demonstrating that the JBU module captures more visual details, enabling precise semantic discrimination.
The choice of grid sizes influences the performance of MLLMs. We explore the impact of RoI-align grid size on the efficiency and performance of LLaVA-UHD v2. Specifically, we RoI-align a region-level feature pyramid from the inverse feature pyramid within a window set, ensuring that higher-resolution feature maps retain finer-grained pooling grids, like [96]. However, this approach, rather than improving multi-scale feature integration, significantly degrades performance and inference efficiency, as demonstrated in Table 4. Compared to fixed grids, selecting a proper pooling grid (defined in Eq.5), offers better performance and efficiency, thanks to a more approximate aspect ratio to the original image.
4.5 Visualization Analysis
Case study. In Fig. 6, we visualize the performance of well-known MLLMs on high-resolution, complex perception tasks. This kind of task requires MLLMs to well fuse both visual details and high-level semantics to accurately identify fine-grained targets ($e.g.$, OCR, colors) during the procedure of complex semantic perception ($e.g.$, semantic relation and visual behavior). It is evident that LLaVA-UHD v2 correctly recognizes the tree planter in the newspaper photo and associates it with the name within the dense image caption (Case A). We also can see that LLaVA-UHD v2 captures the player who raises his hands and reads the “number 3” on his clothes (Case B). In contrast, LLaVA-Next overlooks the name information within dense texts (Case A) and hallucinates on the player number (Case B). Mini-Gemini fails to extract the true name (Case A) and also hallucinates (Case B). Additionally, GPT-4V shows limitations in referencing the information in the newspaper (Case A) and falsely recognizes “number 24” due to wrong fine-grained action perception (Case B). Appendix B.2.2 provides further cases for a more comprehensive comparison across various task scenarios.
Figure 7: Activation response of specific textual tokens to different visual feature levels. Red circles highlight the obvious difference between levels. (Best viewed in color and zoomed-in)
Semantic activation cross feature scale. In Fig. 7, we demonstrate the activation responses of specific textual prompts in the language model to the inverse feature pyramid. As shown, OCR-like textual tokens yield finer-grained and more accurate activations at higher feature levels, facilitating accurate scene text recognition (first row). For object-level semantics, higher feature levels enhance edge detail activations, enabling more precise semantic localization(second row). Collectively, the feature pyramid offers a more exhaustive set of visual semantics with rich granularity, effectively supporting nuanced language decoding.
5 Conclusion
The proposed hierarchical window transformer, central to our LLaVA-UHD v2, effectively addresses the limitations of conventional ViT-based MLLMs by capturing varying visual granularity essential for precise language generation. The Hiwin transformer adeptly constructs an inverse feature pyramid for enriched semantics representation, which is subsequently condensed into a compact set of visual tokens through our hierarchical window attention. This process enhances nuanced visual-linguistic alignment as well as facilitates efficient visual prompting for the LLM. LLaVA-UHD v2 has demonstrated substantial gains over the baseline method across a range of MLLM benchmarks, demonstrating its capability in MLLM tasks that demand both fine-grained and high-level semantics. Furthermore, the Hiwin transformer offers versatility, presenting potential adaptability across diverse ViT-based MLLM architectures.
Appendix A Implementation Details
We provide details of the JBU module, experimental setting on visual tasks, and the dataset composition in the supervised fine-tuning phase of MLLM.
A.1 JBU module
Up-sampling kernel. As presented in Eq. 2 of the main paper, the kernel of the JBU module relies on two weights $D_{dist}$ and $D_{sim}$. Spatial distance decay $D_{dist}$ is defined as
$\begin{split}D_{dist}=\mathrm{exp}\left(-\frac{\lVert(x,y)-(x^{\prime},y^{\prime})\lVert_{2}^{2}}{2\sigma_{dist}^{2}}\right)\end{split}$ (8)
to represent the Euclidean distance relations between adjacent pixels, where $\sigma_{dist}$ denotes a learnable width. And pixel similarity weight $D_{sim}$ is determined as
$\begin{split}D_{sim}=\underset{(x^{\prime},y^{\prime})\in U}{\mathrm{softmax}}\left(\frac{\Theta^{l+1}(\mathcal{I}^{l+1}[x,y])\cdot\Theta^{l+1}(\mathcal{I}^{l+1}[x^{\prime},y^{\prime}])}{\sigma_{sim}^{2}}\right),\end{split}$ (9)
where $\sigma_{sim}$ is a learnable temperature factor to modulate the distribution of similarity scores and $U=7$
Learnable down-sampler. We detail the implementation of the learnable down-sampler defined in Eq. 4 of the main paper. Compared to a simple convolutional layer with a stride of 2, we apply an attention-based down-sampler following  [26] in each feature level. Specifically, a 1$\times$1 convolution layer is first applied to the feature maps of $(l+1)$-th level to extract a saliency map, followed by combining it with a modified fully-connected layer to normalize the features in the local neighborhood $V$. We summarize the above operation as a network $f(\cdot)$ with trainable parameters $\Omega^{l+1}$. As a result, the feature pixel at the location of $[x,y]$ on down-sampled feature maps (of $l$-th level) is formally defined as
$\mathrm{Down}(\mathcal{F}^{l+1};\Omega^{l+1})[{x,y}]=\\ \text{softmax}\left(f(\mathcal{F}^{l+1}[V_{x,y}];\Omega^{l+1})\right)\cdot\mathcal{F}^{l+1}[V_{x,y}],$
where $V_{x,y}$ denotes the local neighborhood in the high-resolution feature maps. Note that, before performing Eq. 10, we experimentally up-sample the $\mathcal{F}^{l+1}$ to the size of the original image using bilinear interpolation. And we set $V_{x,y}=14$, aligning with the patch size of CLIP-ViT, to simulate the feature extraction of ViT.
Multi-view hierarchical reconstruction loss. The objective of the JBU module is to construct high-resolution features by observing multiple different “views” of low-resolution features like NeRF [79] that generates 3D scene representation from multi-view 2D images. Given the input image $\mathcal{I}$ and $l$-th level of feature maps $\mathcal{F}^{l}$, we detail the Eq. 4 of the main paper as
$\mathcal{L}=\frac{1}{2|\mathcal{T}|}\sum_{l=1}^{2}\sum_{t\in\mathcal{T}}\bigg{(}\frac{1}{{{u}}^{2}}\lVert\mathcal{F}^{0}\left(t\left(\mathcal{I}\right)\right)-\mathrm{Down}\left(t\left(\mathcal{F}^{l}\right);\Omega^{l}\right)\rVert_{2}^{2}\\ +\log({u})\bigg{)},$
where $t$ is the image jitter operation sampled from a collection of transformation $\mathcal{T}$ such as pads, zooms, crops, flips, and $etc$. We experimentally set $|\mathcal{T}|=2$. The semantic uncertainty $u=\mathcal{N}(\mathcal{F}^{0}(t(\mathcal{I})))$, as introduced by [33], is parameterized by a linear network $\mathcal{N}$. This term evaluates the certainty of features for up-sampling operations, as such processes may risk losing critical semantic information, which is vital for preserving the integrity of the representation.
Data Size Response formatting prompts
LLaVA [63] 158K
ShareGPT [90] 40K
VQAv2 [29] 83K Answer the question using a single word or phrase.
GQA [38] 72K
OKVQA [75] 9K
OCRVQA [82] 80K
DocVQA [95] 15K
ChartQA [76] 20K
A-OKVQA [88] 66K Answer directly with the option’s letter from the given choices.
DVQA [41] 20K
TextCaps [92] 22K Provide a one-sentence caption for the provided image.
ShareGPT4V [18] 55K
AI2D [43] 3K
LAION-GPT4V [3] 11K
SythDog-EN [46] 40K
LRV-Instruct [61] 30K
RefCOCO 48K
[42, 74] Provide a short description for this region. (for Region Caption)
VG [48] 86K Provide the bounding box coordinate of the region this sentence describes. (for Referring Expression Comprehension)
Total 858K
Table 5: Detailed composition of our 858k-mixed dataset.
A.2 Experimental setting on visual tasks
After training the JBU module in Stage 0, we applied it on the top of the visual backbone to assess its effectiveness. Unless explicitly stated otherwise, both naive bilinear interpolation and the JBU module are employed to perform feature up-sampling on low-resolution (i.e., 24$\times$24) feature maps which are extracted from the second-to-last layer of CLIP-ViT [85].
[…]
Table 6: Comparison of different choices of feature level on performance and efficiency. “HS.”: hierarchical supervision. ESTVQA [101] is a VQA benchmark focusing on scene text recognition.
Efficiency Performance
Level Period(h) Memory(G) Average GQA SQA REC VQAC VQAT ESTVQA MMEP
0,2 27.7 41.9 63.4 63.9 69.5 71.5 60.5 66.5 40.6 71.0
0,1,2 28.0 41.9 63.7 63.8 70.2 71.8 60.5 66.9 40.8 72.1
0,1,2,3 45.6 53.0 63.8 64.4 69.3 72.6 60.7 66.4 41.6 71.4
0,1,2,3 (w/o HS.) 45.6 52.6 62.4 63.6 69.8 67.1 57.8 66.5 39.9 72.0
Figure 8: Qualitative comparison on high-resolution dense perception task which requires the capabilities of fine-grained details perception.
Figure 9: Qualitative comparison on high-resolution fine-grained perception task which requires robust fine-grained visual texture perception capabilities.
Figure 10: Qualitative comparison on high-resolution spatial perception which necessitates the capabilities of high-level spatial contexts.
Figure 11: PCA visualization of the up-sampled features by JBU module on nature scene. With hierarchical supervision, the high-resolution features (8×8$\times$8 ×) could clearly depict object boundary and text appearance. (Best viewed in color and zoomed in)
Figure 12: PCA visualization of the up-sampled features by JBU module on OCR scene. With hierarchical supervision, the high-resolution features (8×8$\times$8 ×) could clearly depict object boundary and text appearance. (Best viewed in color and zoomed in)
A.3 Supervised fine-tuning dataset
As illustrated in Table 5, we detail the proposed 858k-mixed dataset in the supervised fine-tuning phase of MLLM.
Appendix B Analysis
We analyze the behaviors of LLaVA-UHD v2 through qualitative and quantitative experiments.
B.1 Quantitative experiment
Level choice. As shown in Table 6, the introduction of higher level (higher resolution) feature maps results in consistent enhancement of the average performance. However, incorporation of even higher resolution features, such as level-3 feature maps ($i.e.$, 8×8$\times$8 × resolution than level-0), yields marginal benefits while substantially increasing the training cost. If we abandon the hierarchical supervision strategy, performance suffers, which we attribute to the detailed degeneration issue discussed in Sec.3.2 of the main paper. In this case, the JBU module encounters significant challenges in effectively integrating high-frequency information into CLIP-ViT features.
B.2 Qualitative experiment
B.2.1 Enhanced high-resolution features.
As discussed in Sec.3.2 of the main paper, we performed a qualitative visualization to assess the impact of hierarchical supervision, presented in Fig.11 and Fig.12, Note that all the high-resolution features are 8×8$\times$8 × resolution than CLIP-ViT features. While bilinear interpolation increases the nominal resolution of features, it fails to enhance the fidelity of image detail representation. In comparison, a naive JBU module captures finer details but retains a degree of blurriness. With hierarchical supervision, the degradation of high-resolution features is markedly reduced, resulting in a more accurate and refined representation of visual details.
B.2.2 Case studies.
We add more cases to analyze the behavior of our LLaVA-UHD v2. The capabilities are summarized as four aspects as follows.
Dense Perception. In Fig. 8, we visualize the performance of well-known MLLMs on dense perception tasks. Dense perception tasks require models to possess highly robust fine-grained perception capabilities to distinguish object boundaries within a large number of densely packed similar objects, thereby accurately locating the target and its boundaries to identify the target precisely.
It is evident that LLaVA-UHD v2 and GPT-4V accurately identify the beginning time of the television program ‘H20 X5 Mop’ (case A), the performing date of the show ‘The 3 Mile’ (case B), the duration of whole workouts (case C), and the prize of stoli (case D), indicating highly robust fine-grained perception capabilities provided by our visual pyramid representation. In comparison, other models either fail to precisely locate the target (LLaVA-Next) or cannot distinguish the target from similar adjacent objects, limited in accurately completing dense OCR tasks (Mini-Gemini).
Fine-grained Perception. In Fig. 9, we visualized the performance of well-known MLLM on fine-grained perception tasks. These tasks require models to have robust fine-grained perception capabilities to detect the textures of small or blurry targets, thereby accurately locating and identifying small targets.
Cases C indicate that LLaVA-UHD v2 accurately identified the small green light, and the tiny number of duration time associated with green light, demonstrating that the introduction of high-frequency information in hierarchical features can handle small, blurry targets effectively. In contrast, other models, can’t find the small green light, or fail to accurately perform OCR tasks due to the text being too small or blurry (e.g., GPT-4V, LLaVA-Next, Mini-Gemini).
This capability is further demonstrated in cases A, B and D, where both LLaVA-UHD v2 and GPT-4V accurately identified the tiny number on the basketball jersey (case A), the blurry number on the very end of the bus (case B), and the time on the phone (case D), while LLaVA-Next and Mini-Gemini exhibited limitations.
Spatial Perception. In Fig. 10, we visualized the performance of well-known MLLM on spatial perception tasks. Spatial perception tasks require models to have robust high-semantic perception capabilities to discern the spatial relationships between different objects.
It is evident that LLaVA-UHD v2 and GPT4V perceive the spatial relative positions between different objects, to accurately identify the number above the right tire on the car (case A), the number of the player in the middle (case B), the number of the coin on the middle right (case C), the lowest ml (case D). This accuracy is attributed to our high-resolution visual pyramid representation, which allows the perfect integration of features of varying semantics and spatial information across different levels. In contrast, other models, such as LLaVA-Next and Mini-Gemini, fail to accurately perceive these relative spatial positions.